OpenStudy (anonymous):

The particle of dust has a mass of 7.53×10^-10 kilograms. Find the weight of 5 billion dust particles. Plzzz help !!

OpenStudy (anonymous):

\[5.0 \times 10^{9}particles \times \frac{ 7.53 \times 10^{-10}kg }{ 1 particle } \]

OpenStudy (anonymous):

That will give you weight in kg. If you need it in something else, you need to convert kg to that something else

OpenStudy (anonymous):

@Larryboxaplenty so I'll be dividing 7.53×10^-10 by 5,000,000,000?

OpenStudy (anonymous):

 You multiply the mass of one particle by 5 billion. The one is there so show that the unit "particles" cancel